A client was prescribed 10 mg of a medication for a week. After a week, the health care provider reduced the medication to seven (7 mg). What was the percentage of decrease in medication
Answer #1

Ans) Initial dose: 10 mg

Decreased the dose to 7 mg so 10-7 = 3

Percentage decrease= 3/10 = 0.3%